A system absorbs 41 J of heat. And the same time the system does 6 J of work on the surroundings. What is the ∆U?

A system releases 35 J of heat. And the same time 2 J of work is done on the system by the surroundings. What is the ∆U?

A system releases 16 J of heat. And the same time 27 J of work is done on the system by the surroundings. What is the ∆U?

A system absorbs 76 J of heat. And the same time the system does 21 J of work on the surroundings. What is the ∆U?
Answer #1

1)
Q = +41 J
W = -6 J

Use:
∆U = Q + W
= 41 J - 6 J
= 35 J
Answer: 35 J

2)
Q = -35 J
W = +2 J

Use:
∆U = Q + W
= -35 J + 2 J
= -33 J
Answer: -33 J

3)
Q = -16 J
W = +27 J

Use:
∆U = Q + W
= -16 J + 27 J
= 11 J
Answer: 11 J

4)
Q = +76 J
W = -21 J

Use:
∆U = Q + W
= 76 J - 21 J
= 55 J
Answer: 55 J